a game spinner has regions that are numbered 1 through 8. If the spinner is used twice, what is the probablility that the first number is a 3 and the second is an even number?
Answer by htmentor(1343)

The total probability is the product of the individual probabilities.
The probability of getting a particular number, in this case a 3, out of the eight possible numbers is:
P(1st number is 3) = 1/8
Out of the 8 numbers, 4 are even. Therefore:
P(even number) = 4/8 = 1/2
So the total probability is:
P(3 and even) = 1/8*1/2 = 1/16